OpenStudy (anonymous):

John and Barry are pulling on a fence post with ropes while their friend Pat uses a chain-saw to sheer it at ground-level. John applies a force of 75.0N [40*] on one rope while Barry applies a force of 85.0N [340*] on the other rope. What is the net force on the post? Show your work.

OpenStudy (therealmeeeee):

ummmmm

OpenStudy (anonymous):

haha Ya i can not help you. sorry

OpenStudy (anonymous):

@waterpologoalie Fn = 75N[40o] + 85N[340o] X = 75*cos40 + 85*cos340 = 137.3 N. Y = 75*sin40 + 85*sin340 = 19.14 N. tan A = Y/X = 19.14/137.3 = 0.13940 A = 7.94o Fn=X/cosA = 137.3/cos7.94=138.6 N[7.94o]